OpenStudy (anonymous):

Fan and medal! Can someone check my work please? The first challenge to complete is the Penny Problem. The radio station is giving the winner of this challenge a prize pack that includes tickets to see his or her favorite band in concert. To start off the challenge, the radio station has placed pennies in a cylindrical glass jar. Each penny is 0.75 inches in diameter and 0.061 inches thick. If the cylindrical glass jar containing the pennies has a diameter of 6 inches and a height of 11.5 inches, how many pennies can fit inside the jar? You must show all work to receive credit.

OpenStudy (anonymous):

Each Penny is 0.75 in diameter, and 0.065 inches thick. Cylindrical Glass jar is 6 inches in diameter, and 11.5 inches in height. Volume of cylindrical glass V = 3.14r^2 h V = 3.14 (3)^2 (11.5) V = 3.14 (9) (11.5) V = 3.14 (103.5) V = 324.99 inches. Volume of Pennies V = 3.14r ^2 h V = 3.14 (0.375)^2 (0.061) V = 0.4415625 ( 0.061) V = 0.026935312 Now divide the volume of the glass by the volume of the pennies. And round it to the nearest hundredths. 324.99 / 0.026935312 = 12065.57399 rounded to the nearest hundredths 12065.57

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@thomaster